# Unleashing the Power of Large Language Models for Group POI Recommendations

### v1

@@ 3.3. Self-supervised Leaning Task
3.4. The Optimization of LLMGPR
The optimization workflow of LLMGPR is presented in Algorithm 1. At the very beginning (line 1), we regard Llama3-8b as the base large language model $\Theta$, which can be replaced by any other large language model if sufficient computational resources are available. After that, we initialize all trainable weights including the check-in sequencing adapter $\Theta^{{}^{\prime}}_{s}$, the aggregation adapter $\Theta^{{}^{\prime}}_{a}$, the POI embedding matrix $\textbf{E}_{poi}$, and the purpose embedding matrix $\textbf{E}_{pur}$ (line 2). Then, we obtain the sequence-specific LLM $\Theta_{s}$ and the aggregation-specific LLM $\Theta_{a}$ (line 3-4). On this basis, we first pre-train the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$ based on the purpose prediction loss $\mathcal{L}_{pur}(y^{\prime}_{p},y_{p})$ using all short sequences and their corresponding purpose labels (lines 5-10). The sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$ are further optimized by the POI prediction loss $\mathcal{L}_{seq}(y^{\prime},y)$ (lines 11-16). To address the group check-in data sparsity and obtain a better task-specific LLM for learning the representation of check-in sequences, all check-in sequences are utilized including group and all individual check-in sequences.
Subsequently, the aggregation adapter $\Theta^{{}^{\prime}}_{a}$ is trained iteratively with group check-in sequences and corresponding individual check-in sequences of each group member (lines 17-29). In detail, for each group check-in sequence, the group representation is originally obtained by applying the sequence-specific LLM $\Theta_{s}$ to the group check-in sequence (line 19). After that, individual representations for group members are obtained by applying $\Theta_{s}$ to their respective check-in sequences (lines 20-23). These representations are then fed into the aggregation-specific LLM $\Theta_{a}$ (line 24), while the aggregated representation is utilized to enhance the group representation (line 25) and generate the recommendation results (line 26). Finally, the aggregation adapter is updated using the POI prediction loss $\mathcal{L}_{seq}(y^{\prime},y)$ (line 27).
4. Experiments
In this section, we perform comprehensive experiments with two real-world datasets to evaluate the effectiveness of the proposed LLMGPR. Specifically, our investigation seeks to address the following research questions:
RQ1: How effective is LLMGPR for group POI recommendations?
RQ3: Is each key component in LLMGPR effective?
RQ4: Does LLMGPR work well for cold-start recommendations?
RQ5: What is the influence of LLMGPR’s key hyperparameters?
4.1. Datasets and Evaluation Protocols
We adopt two real-world datasets including Foursquare (Chen et al., 2020a) and Weeplace (Liu et al., 2013) to evaluate the proposed LLMGPR, where both datasets include check-in histories in the city of New York. Since both datasets do not originally contain group information, we follow the widely adopted procedure (Chen et al., 2022) to construct group interactions by finding overlaps in check-in times and social relations. That is, we assume if a set of users who are connected on the social network visit the same venue at the same time, then they are regarded as members of a group, and the corresponding activities are group activities. The use of explicit social connections and spatiotemporal tags ensures the quality of discovered user groups. Table 5 summarizes the statistics of the two datasets.
For evaluation, we adopt the leave-one-out protocol which is widely used in previous works (Wang et al., 2018, 2019; Zheng et al., 2016). That is, for each of the check-in sequences, the last check-in POI is for testing, the second last POI is for validation, and all others are for training. In addition, the maximum sequence length is set to 200. For each ground truth check-in POI, unlike the E-commerce setting that ranks all products (Krichene and Rendle, 2020), we only pair it with 500 unvisited and nearest POIs within the same region as the candidates for ranking. The rationale is, different from e-commerce products (Krichene and Rendle, 2020; Qu et al., 2021), in the scenario of POI recommendations that are location-sensitive, users seldom travel between two POIs consecutively that are far away from each other (Long et al., 2023; Yuan et al., 2023; Li et al., 2021). On this basis, we leverage two ranking metrics, namely Hit Ratio at Rank $k$ (HR@$k$) and Normalized Discounted Cumulative Gain at Rank $k$ (NDCG@$k$) (Weimer et al., 2007) where HR@$k$ only measures the times that the ground truth is present on the top-$k$ list, while NDCG@$k$ cares whether the ground truth can be ranked as highly as possible.
4.2. Baselines
To evaluate the effectiveness of LLMGPR for group POI recommendations, we first compare it with existing group recommenders: MF-AVG (Vinh Tran et al., 2019), AGREE (Cao et al., 2018), GroupIM (Cao et al., 2018), CubeRec (Chen et al., 2022), HHGR (Chen et al., 2022) and MICL (Li and Zhu, 2024). For a fair comparison, we regard check-in sequence representations and POI representations generated by the well-trained LLM as initial representations of groups, users, and POIs for the above baselines. To validate the effectiveness of LLMGPR in learning check-in sequence representations, we also compare it with SOTA POI recommenders for individuals: GeoMF (Lian et al., 2014), STAN (Luo et al., 2021), DRAN (Wang et al., 2022), Diff-POI (Qin et al., 2023), META ID (Huang et al., 2024), LLMMove (Feng et al., 2024) and LLM4POI (Li et al., 2024). The details of those baselines are provided in Appendix A.4.
Foursquare Weeplace
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
MF-AVG 0.2560 0.1480 0.3666 0.2143 0.2597 0.1501 0.3896 0.2320
AGREE 0.2825 0.1590 0.4003 0.2327 0.2713 0.1614 0.4226 0.2542
GroupIM 0.2917 0.1663 0.4060 0.2467 0.2852 0.1678 0.4459 0.2511
CubeRec 0.3012 0.1683 0.4170 0.2509 0.2926 0.1714 0.4567 0.2584
HHGR 0.3086 0.1795 0.4391 0.2585 0.3033 0.1825 0.4637 0.2710
MICL 0.3101 0.1797 0.4516 0.2653 0.3121 0.1834 0.4693 0.2766
LLMGPR 0.3198 0.1926 0.4875 0.2874 0.3217 0.1917 0.4978 0.2934
4.3. Group Recommendation Effectiveness (RQ1)
The performance comparison among all the group recommenders is summarized in Table 1, revealing several key insights. Firstly, AGREE consistently outperforms MF-AVG on both datasets, due to its attention mechanism that dynamically assigns weights for preference aggregation, thus better capturing the interactions and relationships among group members. However, the use of point embeddings limits its flexibility and resilience in representing groups. To address this, GroupIM and CubeRec introduce hypercube-based group representations, replacing point embeddings with range embeddings, which achieve better recommendation accuracy. Additionally, HHGR and MICL utilize graph neural networks to capture relationships between users, groups, and POIs, resulting in more precise group representations and more accurate recommendations. Despite the advances of these methods, the proposed LLMGPR model demonstrates superior accuracy over all baselines. Specifically, compared to the best-performing baseline, MICL, LLMGPR achieves average improvements of 5.05% on Foursquare and 4.86% on Weeplace. Another notable observation is that all the group recommenders achieve higher recommendation accuracy on Weeplace than on Foursquare, due to its sparser group-POI interactions and a larger set of POIs. Notably, the proposed LLMGPR can still leverage augmented supervision signals to learn quality group representations, achieving the best group recommendation performance even on this challenging dataset.
Foursquare Weeplace
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
GeoMF 0.1051 0.0604 0.1326 0.0892 0.0849 0.0610 0.0958 0.0664
STAN 0.3113 0.1801 0.4654 0.2896 0.3004 0.1761 0.4361 0.2621
DRAN 0.3159 0.1826 0.4816 0.2970 0.3107 0.1807 0.4349 0.2677
Diff-POI 0.3211 0.1935 0.4870 0.2918 0.3106 0.1848 0.4751 0.2827
META ID 0.3058 0.1799 0.4711 0.2915 0.3075 0.1813 0.4338 0.2702
LLMMove 0.3146 0.1892 0.4737 0.2813 0.3062 0.1829 0.4763 0.2763
LLM4POI 0.3241 0.1996 0.4976 0.2970 0.3133 0.1882 0.4898 0.2936
LLMGPR 0.3386 0.2129 0.5101 0.3037 0.3284 0.2027 0.5072 0.3012
To further study the effectiveness of the proposed LLMGPR in learning check-in sequence representation, we also compare LLMGPR with SOTA POI recommenders for individuals, where the results are summarized in Table 2. Our proposed LLMGPR consistently and significantly beats the baselines on both datasets. As the basic POI recommendation approach without considering the sequential effect of user check-in activities, GeoMF undoubtedly has the worst performance on both datasets. STAN outperforms GeoMF by a significant margin since it can leverage spatiotemporal correlations in both consecutive and non-consecutive check-in activities with the attention mechanism. Advancing further, DRAN melds a Graph Neural Network (GNN) with an attention mechanism, leading to more refined POI embeddings and thus outperforming STAN in terms of accuracy. Most notably, Diff-POI, employing the robust generality of the diffusion model, sets a new benchmark for state-of-the-art accuracy in this domain.
Recent works focus on integrating the LLM for recommendations, due to its ability to leverage the semantic information of users and items. META ID uses out-of-vocabulary tokens to characterize users and items for recommendations. As it is not designed for POI recommendations, it cannot integrate and exploit the spatial and temporal information of check-ins, even performing worse than Diff-POI. In addition, LLMMove applies the zero-shot generation ability of the LLM for the next POI recommendations. Without fine-tuning for the specific recommendation task, its performance is limited. Instead, LLM4POI fine-tunes the LLM on the two datasets for the POI recommendation task, and thus, it achieves better performance. Notably, our proposed LLMGPR significantly outperforms the best baseline LLM4POI, showing its superiority in learning check-in sequence representations. This is attributed to the well-trained sequencing adapter and the introduction of a novel self-supervised learning task.
Table 3. Ablation study for LLMGPR.
Foursquare Weeplace
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
LLMGPR 0.3198 0.1926 0.4875 0.2874 0.3217 0.1917 0.4978 0.2934
LLMGPR-FT 0.2781 0.1652 0.4159 0.2355 0.2947 0.1704 0.4312 0.2439
LLMGPR-M 0.2834 0.1688 0.4367 0.2495 0.2912 0.1704 0.4472 0.2592
LLMGPR-ER 0.2921 0.1701 0.4332 0.2524 0.2945 0.1719 0.4513 0.2666
LLMGPR-SSL 0.3119 0.1816 0.4623 0.2781 0.3091 0.1825 0.4732 0.2847
4.5. Ablation Study (RQ3)
In this section, we perform the ablation analysis to evaluate the performance gain from the core components of LLMGPR in terms of the main task: group POI recommendations. Table 3 summarizes the recommendation outcomes for different degraded versions of LLMGPR. Then, we introduce all variants and discuss the effectiveness of corresponding model components.
Using the LLM directly without fine-tuning (LLMGPR-FT). In LLMGPR, we modify and fine-tune the LLM for group POI recommendations. To evaluate the usefulness of those modifications, we apply the original LLM directly for group POI recommendations, where both group-level and member-level check-in sequences are utilized. This variant model has experienced noticeable performance drops, especially on the Foursquare dataset with the higher sparsity of group-item interactions. Hence, including modification and fine-tuning of the LLM is important for LLMGPR to learn representative group representations in the scenario of data sparsity.
Removing the enhancement of group representation with aggregated member preferences (LLMGPR-ER). A crucial component of LLMGPR is enhancing group representation by aggregated member preference as described in Section 3.2. To test its effectiveness, we remove it and only utilize the group check sequence for final recommendations. As a result, there is a significant decrease in recommendation accuracy for both datasets. This reflects that, even with the well-trained LLM, the sparse group-item interactions significantly harm the accuracy, and our proposed aggregation strategy can effectively address this issue.
Removing self-supervised learning (LLMGPR-SSL). This variant disables the self-supervised loss when pretraining the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$, as shown in Section 3.3. As LLMGPR can no longer leverage the purposes of short sequences as a supervision signal to counteract the group-level data sparsity, it suffers from inferior recommendation accuracy on both datasets. Therefore, the novel self-supervised learning scheme showcases its strong contribution to enriching group-level information.
4.6. Effect of Cold-start Recommendations (RQ4)
To demonstrate LLMGPR’s ability in cold-start recommendations, we randomly selected $200$ group-level check-in sequences from each dataset. These sequences were excluded from the training process and each had fewer than 10 check-ins. We then recorded the average recommendation accuracy (HR@10) for these sequences across all group POI recommenders, including LLMGPR, as shown in Table 4. LLMGPR’s cold-start accuracy was notably higher than all baselines, highlighting its effectiveness in providing recommendations for new groups with limited historical check-in data.
Table 4. Cold-start recommendation performance (HR@10).
Foursquare Weeplace
MF-AVG 0.3148 0.3203
AGREE 0.3605 0.3537
GroupIM 0.3698 0.3849
CubeRec 0.3811 0.3925
HHGR 0.4178 0.4274
MICL 0.4157 0.4192
LLMGPR 0.4689 0.4663
4.7. Hyperparameter Sensitivity (RQ5)
Impact of $r$. Recommendation accuracy is recorded for $\lambda\in\{4,8,16,32,64\}$. Usually, the recommendation accuracy benefits from higher $r$. However, the accuracy stabilizes once $r$ exceeds $16$, indicating that LLMGPR is capable of delivering high-performance recommendations without requiring additional resources to train larger adapters.
Figure 6. Hyperparameter sensitivity.
5. Related work
5.1. POI Recommendations
To help people discover attractive places by analyzing user-POI interactions, early works on discovering attractive places through user-POI interactions primarily focused on matrix factorization (Lian et al., 2014) and Markov chains (Cheng et al., 2013; Zheng et al., 2016). The advent of recurrent neural network (RNN) models marked a significant advancement, as they demonstrated the ability to capture spatiotemporal dynamics in POI sequences (Chen et al., 2020b; Jie et al., 2018; Chen et al., 2020c; Yin et al., 2015; Li et al., 2021). Additionally, models utilizing attentive neural networks (Luo et al., 2021; Yang et al., 2022; Chen et al., 2019; Yin et al., 2015) adopted self-attention mechanisms to analyze the spatiotemporal context of sequential check-ins. Graph neural network (GNN)-based models (Li et al., 2021; Wang et al., 2022; Gao et al., 2023a, b) further advanced the field by integrating graph-augmented POI sequences, leveraging collaborative signals from semantically similar POIs to reveal sequential trends and outperform RNN-based approaches in accuracy. Diff-POI (Qin et al., 2023) sets a new standard by using diffusion models to achieve cutting-edge accuracy. Recent research focuses on integrating Large Language Models (LLMs) for POI recommendations. LLMMove (Feng et al., 2024) employs zero-shot generation for next POI recommendations, while LLM4POI (Li et al., 2024) fine-tunes LLMs on related datasets to exploit commonsense knowledge. Compared to these approaches, LLMGPR demonstrates significant advantages in learning sequence representations and addressing group-level data sparsity by optimizing adapters and the POI embedding matrix.
5.2. Group POI Recommendations
Group POI recommenders typically involve two main phases: learning group representations from group-level check-in sequences and enhancing these representations using aggregated member representations derived from individual check-in sequences. Existing methods for preference aggregation range from simple averaging (Vinh Tran et al., 2019), which treats all members equally, to weighted averaging (Sojahrood and Taleai, 2021), which assigns weights based on external factors but struggles with determining appropriate values. More sophisticated approaches, such as attentive neural networks (Cao et al., 2018; Guo et al., 2020), hypercube-based embeddings (Sankar et al., 2020; Chen et al., 2022), and graph convolution networks (Zhang et al., 2021; Li and Zhu, 2024), aim to capture multifaceted preferences and complex relationships among users, groups, and POIs. However, these models often face challenges like overfitting and lack of scalability. In contrast, LLMGPR, a modified and well-trained LLM, can capture complex patterns and adapt dynamically to individual preferences, offering improved performance and adaptability for group POI recommendations.
5.3. LLM-based Recommendations
Motivated by the notable achievements of LLMS, the integration of LLMs into recommendation systems has grown significantly. Early works (Feng et al., 2024; Hou et al., 2024) primarily focused on designing prompts to utilize LLMS directly with zero-shot generalization. However, these approaches often struggled with the lack of specificity and domain adaptation, which led to the development of prompt tuning techniques (Zhai et al., 2023; Wu et al., 2024). Prompt tuning involved adjusting the prompts to better suit the recommendation task, offering improved performance but still facing limitations in handling complex recommendation scenarios. This led to the exploration of fine-tuning LLMs (Lin et al., 2024; Bao et al., 2023), where models were adapted to the recommendation tasks through task-specific training, significantly enhancing accuracy but lacking unique indicates of item IDs. Hence, the construction of IDs was introduced (Huang et al., 2024; Yu et al., 2024), creating embeddings or representations that could capture nuanced user preferences and item characteristics, thus offering a more efficient and scalable solution to recommendation problems. LLMGPR introduces and optimizes additional weights, including POI embeddings, sequencing adapters, and aggregation adapters, to modify the LLM for group-level recommendations. These modifications aim to learn sequence representations and aggregate members’ preferences more effectively.
6. Conclusion
In this paper, we introduced LLMGPR, a novel large language model-based framework designed to address the challenges of group POI recommendation in Location-Based Social Networks. Our approach effectively learns sequence representations and dynamically aggregates group members’ preferences, overcoming the limitations of existing methods that struggle with data sparsity and inadequate preference aggregation. By leveraging the capabilities of large language models and introducing additional POI tokens, and adapters for sequencing and aggregation tasks, LLMGPR generates dense, holistic representations that improve recommendation accuracy. Additionally, a self-supervised learning signal enhances performance and generalizability across diverse groups. Our experiments demonstrate significant improvements over existing methods, highlighting the potential of LLMGPR to revolutionize group POI recommendations and better satisfy the diverse preferences of group members.
Appendix A Appendix
A.1. Short sequence purple labels
All short sequence labels for self-supervised learning in Section 3.3 are:
  • Business Trip: This label captures check-ins related to work travel, such as meetings, conferences, or site visits. Business trips are often brief but involve specific destinations like hotels, airports, and office buildings, making them important for identifying professional user behaviors.
  • Work Commute: This label covers routine, daily commutes to and from work. It captures repetitive patterns in user check-ins tied to their professional lives, providing useful insights into regular, predictable movements between home and work locations.
  • Tourism: This category includes check-ins at tourist attractions, such as landmarks, museums, or vacation destinations. Tourism-related trips are typically short in duration but involve diverse POIs, making it a crucial label for understanding users’ travel behavior.
  • Short Vacation: This label captures brief leisure trips that are not necessarily tied to family or friends but are characterized by relaxation and exploration over a few days. Such vacations may include weekend getaways, short stays at resorts, or visits to nearby tourist destinations. Unlike long trips, short vacations typically involve fewer check-ins but reflect an intention to explore new locations within a limited time, providing valuable insight into recreational user behavior.
  • Cultural & Entertainment: This label covers activities like visiting theaters, concerts, galleries, or other entertainment venues. Cultural outings often represent unique or occasional check-ins, distinct from routine patterns, and are driven by leisure and enjoyment.
  • Shopping & Dining: This category encompasses visits to shopping centers, retail stores, and restaurants. These activities are typically short-term, common in urban exploration, and represent essential behaviors in users’ leisure and daily lives.
  • Social Events: This label covers participation in larger gatherings like weddings, parties, or festivals. Social events drive check-ins at specific venues and are typically tied to one-off, significant occasions.
  • Fitness & Wellness: This category includes gym visits, outdoor runs, or wellness centers. These check-ins often happen regularly within short time spans, reflecting health-focused user habits.
  • Daily Commute: This label captures non-work-related daily travel, such as trips to grocery stores, schools, or other essential locations. These are habitual, short-duration check-ins critical for understanding routine user behavior outside of work.
  • Medical & Health Visits: This label covers check-ins at hospitals, clinics, or pharmacies. These visits are generally short-term and occur due to specific health needs, adding an important dimension to occasional user behavior.
  • Education & Training: This category includes check-ins related to learning environments, such as schools, universities, or training centers. These trips often involve short-term attendance but are crucial for understanding education-driven travel.
These labels are designed to reflect the most common and distinct purposes for short-term trips or sequences of check-ins. Each category is tied to a specific motivation or context that drives user mobility, ranging from daily routines to more occasional events. Together, they comprehensively capture the various reasons behind short-duration check-ins, ensuring that each sequence can be classified meaningfully and relevantly for POI prediction.
A.2. Algorithm 1: The optimization of LLMGPR
@@
16:      Take a gradient step w.r.t $\mathcal{L}_{seq}(y^{\prime},y)$ to update $\Theta^{{}^{\prime}}_{s}$, $\textbf{E}_{poi}$;
@@
30:      Take a gradient step w.r.t $\mathcal{L}_{seq}(y^{\prime},y)$ to update $\Theta^{{}^{\prime}}_{a}$;
@@
A.3. Dataset statistics.
@@
Foursquare Weeplace
#users 6,078 3,795
#groups 1,557 867
#POIs 63,445 35,911
#categories 436 625
#user check-ins 923,856 519,915
#group check-ins 10,899 10,404
#check-ins per user 152 137
#check-ins per group 7 12
#users per group 3.67 4.58
A.4. Baselines

### v2

@@ 3.3. Addressing Data Sparsity with Self-supervised Leaning
3.4. Optimization
@@
16:   Take a gradient step w.r.t $\mathcal{L}_{poi}(y^{\prime},y)$ to update $\Theta^{{}^{\prime}}_{s}$, $\textbf{E}_{poi}$;
@@
30:   Take a gradient step w.r.t $\mathcal{L}_{poi}(y^{\prime},y)$ to update $\Theta^{{}^{\prime}}_{a}$;
@@
The optimization workflow of LLMGPR is presented in Algorithm 1. At the very beginning (line 1), we regard Llama3-8b as the base large language model $\Theta$, which can be replaced by any other large language model if sufficient computational resources are available. After that, we initialize all trainable weights including the check-in sequencing adapter $\Theta^{{}^{\prime}}_{s}$, the aggregation adapter $\Theta^{{}^{\prime}}_{a}$, the POI embedding matrix $\textbf{E}_{poi}$, and the purpose embedding matrix $\textbf{E}_{pur}$ (line 2). Then, we obtain the sequence-specific LLM $\Theta_{s}$ and the aggregation-specific LLM $\Theta_{a}$ (line 3-4). On this basis, we first pre-train the sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$ based on the purpose prediction loss $\mathcal{L}_{pur}(y^{\prime}_{p},y_{p})$ using all short sequences and their corresponding purpose labels (lines 5-10). The sequencing adapter $\Theta^{{}^{\prime}}_{s}$ and the POI embedding matrix $\textbf{E}_{poi}$ are further optimized by the POI prediction loss $\mathcal{L}_{poi}(y^{\prime},y)$ (lines 11-16). To address the group check-in data sparsity and obtain a better task-specific LLM for learning the representation of check-in sequences, all check-in sequences are utilized including group and all individual check-in sequences.
Subsequently, the aggregation adapter $\Theta^{{}^{\prime}}_{a}$ is trained iteratively with group check-in sequences and corresponding individual check-in sequences of each group member (lines 17-29). In detail, for each group check-in sequence, the group representation is originally obtained by applying the sequence-specific LLM $\Theta_{s}$ to the group check-in sequence (line 19). After that, individual representations for group members are obtained by applying $\Theta_{s}$ to their respective check-in sequences (lines 20-23). These representations are then fed into the aggregation-specific LLM $\Theta_{a}$ (line 24), while the aggregated representation is utilized to enhance the group representation (line 25) and generate the recommendation results (line 26). Finally, the aggregation adapter is updated using the POI prediction loss $\mathcal{L}_{poi}(y^{\prime},y)$ (line 27).
4. Experiments
In this section, we perform comprehensive experiments with three real-world datasets to evaluate the effectiveness of the proposed LLMGPR. Specifically, our investigation seeks to address the following research questions:
RQ1: How effective is LLMGPR for group POI recommendations?
RQ3: Is each key component in LLMGPR effective?
RQ4: Does LLMGPR work well for cold-start recommendations?
RQ5: What is the influence of LLMGPR’s key hyperparameters?
@@ 4. Experiments
Foursquare Weeplace Gowalla
#users 7,507 4,560 31,751
#groups 1,715 923 2186
#POIs 80,962 44,194 81,123
#categories 436 625 537
#user check-ins 1,214,631 623,654 862,502
#group check-ins 12,594 11,974 7,738
#check-ins per user 162.80 136.77 27.16
#check-ins per group 7.34 12.97 3.54
#users per group 3.72 4.37 2.95
4.1. Datasets and Evaluation Protocols
We adopt three real-world datasets including Foursquare (Chen et al., 2020), Weeplace (Long et al., 2023b), and Gowalla (Liu et al., 2013) to evaluate the proposed LLMGPR, where all datasets include user’s check-in histories in the cities of New York, Los Angeles, and Chicago. Following (Li et al., 2018; Long et al., 2024a), users and POIs with less than 10 interactions are removed. Since all datasets do not originally contain group information, we follow the widely adopted procedure (Chen et al., 2022) to construct group interactions by finding overlaps in check-in times and social relations. That is, we assume if a set of users who are connected on the social network visit the same venue at the same time, then they are regarded as members of a group, and the corresponding activities are group activities. The use of explicit social connections and spatiotemporal tags ensures the quality of discovered user groups. Table 1 summarizes the statistics of the three datasets.
For evaluation, we adopt the leave-one-out protocol which is widely used in previous works (Zheng et al., 2016; Zhang et al., 2023). That is, for each of the check-in sequences, the last check-in POI is for testing, the second last POI is for validation, and all others are for training. In addition, the maximum sequence length is set to 200. For each ground truth check-in POI, unlike the E-commerce setting that ranks all products (Krichene and Rendle, 2020; Yin et al., 2024), we only pair it with 500 unvisited and nearest POIs within the same region as the candidates for ranking. The rationale is, different from e-commerce products (Krichene and Rendle, 2020; Qu et al., 2021), in the scenario of POI recommendations that are location-sensitive, users seldom travel between two POIs consecutively that are far away from each other (Long et al., 2023a; Li et al., 2021). On this basis, we leverage two ranking metrics, namely Hit Ratio at Rank $k$ (HR@$k$) and Normalized Discounted Cumulative Gain at Rank $k$ (NDCG@$k$) (Weimer et al., 2007) where HR@$k$ only measures the times that the ground truth is present on the top-$k$ list, while NDCG@$k$ cares whether the ground truth can be ranked as highly as possible.
Foursquare Weeplace Gowalla
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
MF-AVG 0.2530 0.1457 0.3577 0.2114 0.2578 0.1476 0.3752 0.2279 0.2050 0.1238 0.3205 0.1875
AGREE 0.2731 0.1573 0.3869 0.2303 0.2612 0.1552 0.4201 0.2459 0.2123 0.1306 0.3486 0.2038
GroupIM 0.2863 0.1639 0.3872 0.2420 0.2848 0.1652 0.4431 0.2498 0.2230 0.1314 0.3595 0.2009
CubeRec 0.2919 0.1638 0.4016 0.2454 0.2826 0.1682 0.4435 0.2582 0.2288 0.1402 0.3656 0.2061
HHGR 0.3057 0.1705 0.4247 0.2584 0.3029 0.1789 0.4409 0.2595 0.2425 0.1495 0.3718 0.2178
MICL 0.3068 0.1711 0.4291 0.2571 0.3043 0.1786 0.4672 0.2676 0.2565 0.1477 0.3808 0.2299
LLMGPR 0.3182 0.1817 0.4767 0.2760 0.3201 0.1856 0.4799 0.2829 0.2739 0.1619 0.4203 0.2512
Foursquare Weeplace Gowalla
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
GeoMF 0.1029 0.0590 0.1563 0.1075 0.0845 0.0599 0.1344 0.0937 0.0706 0.0634 0.1238 0.0955
STAN 0.3112 0.1770 0.4619 0.2759 0.2930 0.1734 0.4155 0.2547 0.2560 0.1520 0.3601 0.2164
DRAN 0.3020 0.1744 0.4769 0.2941 0.3064 0.1726 0.4220 0.2625 0.2579 0.1564 0.3751 0.2335
Diff-POI 0.3119 0.1885 0.4711 0.2787 0.3034 0.1760 0.4681 0.2822 0.2585 0.1543 0.4050 0.2415
META ID 0.2947 0.1721 0.4696 0.2780 0.3064 0.1723 0.4129 0.2615 0.2557 0.1545 0.3728 0.2381
LLMMove 0.3127 0.1799 0.4693 0.2805 0.2958 0.1770 0.4627 0.2657 0.2540 0.1572 0.4044 0.2428
LLM4POI 0.3173 0.1948 0.4776 0.2849 0.3023 0.1874 0.4703 0.2891 0.2699 0.1628 0.4246 0.2492
LLMGPR 0.3233 0.2040 0.4872 0.2995 0.3196 0.1963 0.4867 0.2920 0.2853 0.1958 0.4483 0.2586
4.2. Baselines
To evaluate the effectiveness of LLMGPR for group POI recommendations, we first compare it with existing group recommenders. For a fair comparison, we regard check-in sequence representations and POI representations generated by the well-trained LLM as initial representations of groups, users, and POIs for the above baselines. To validate the effectiveness of LLMGPR in learning check-in sequence representations, we also compare it with SOTA POI recommenders for individuals. All baselines are summarized below: